OpenStudy (anonymous):

Find the product of (x + 9i)^2

OpenStudy (anonymous):

@chetan552

OpenStudy (anonymous):

@solomonzelman please help

OpenStudy (anonymous):

It works the same way as (a+b)^2 just you need to know that i^2=-1

OpenStudy (anonymous):

can you explain step by step please? @Andras

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2 This is the foundation of algebra so I really hope you know that! (x + 9i)^2= x^2+2x9i+(9i)^2 = x^2+18xi +81i^2=x^2+18xi-81

OpenStudy (anonymous):

@Andras thank you so much! can you help me in 2 more? I know how to start it but the foil part is what mostly throws me off plus i started florida virtual late so im a bit lost in what we are doing

OpenStudy (anonymous):

I am going to bed soon. It is just past midnight here. Post a new question and I am sure some1 will help you out.

OpenStudy (anonymous):

Find the product of (x – 2i)^2 Find the product of (x + (3+5i))^2 Thats super sweet of you, thank you so much! @Andras

OpenStudy (anonymous):

(x – 2i)^2 x^2-2x2i+(2i)^2=x^2-4xi-4

OpenStudy (anonymous):

(x + (3+5i))^2 x^2+2x(3+5i)+(3+5i)^2 This has a sub-problem within: (3+5i)^2=9+15i-25 x^2+2x(3+5i)+(3+5i)^2= x^2+6x+10xi+9+15i-25= x^2+6x+10xi+15i-16

OpenStudy (anonymous):

@Andras thank you so much!! :)

OpenStudy (anonymous):

Hope you think it through at least and not just copy-paste for your homework. :0

OpenStudy (anonymous):

@Andras dont worry im reviewing it and noticing how it gets to the final answer!

OpenStudy (anonymous):

Its nothing but (a+b)^2 formula... (a+b)^2= a^2+b^2+2ab hear a=x , b=9i and also dont forget that in the expantion you will see i^2 dont worry about that just repace i^2 with -1 you will get the answer....